# R&D Networks under Heterogeneous Firm Productivities

### v1

@@ Abstract
We introduce heterogeneous R&D productivities into an endogenous R&D network formation model, generalizing the framework in goyal2001r. Heterogeneous productivities endogenously create asymmetric gains for connecting firms: the less productive firm benefits disproportionately, while the more productive firm exerts greater R&D effort and incurs higher costs. For sufficiently large productivity gaps between two firms, the more productive firm experiences reduced profits from being connected to the less productive one. This overturns the benchmark results on pairwise stable networks: for sufficiently large productivity gaps, the complete network becomes unstable, whereas the Positive Assortative (PA) network—where firms cluster by productivity levels—emerges as stable. Simulations show that the PA structure delivers higher welfare than the complete network; nevertheless, welfare under PA formation follows an inverted U-shape in the fraction of high-productivity firms, reflecting crowding-out effects at high fractions. Altogether, a counterintuitive finding emerges: economies with higher average R&D productivity may exhibit lower welfare through (i) the formation of alternative stable R&D network structures or (ii) a crowding-out effect of high-productivity firms. Our findings highlight that productivity-enhancing policies should account for their impact on endogenous R&D alliances and effort, as such endogenous responses may offset or even reverse the intended welfare gains. JEL Classification: D21, D85, O30
@@ 1 Introduction
Innovation is widely acknowledged as a fundamental driver of economic growth in modern societies (romer1990endogenous; jones1995r). A key engine of the innovation process is investment by private firms in R&D (grossman1991quality; aghion1992constructive), a key component of which is collaborative R&D alliances (e.g., as empirically shown in cassiman2002r; belderbos2004cooperative; calero2007research).
@@
In this paper, we study collaborative R&D alliances and their implications for firms’ profits and aggregate welfare in an R&D network formation model, where firms have heterogeneous R&D productivities. A growing theoretical literature, including goyal2001r; goyal2003networks; konig2019r, has studied R&D networks where firms improve their R&D efficiencies by forming collaborations while still competing with each other in the product market. However, these models typically assume that firms are homogeneous in their ability to transform R&D effort into cost-reducing innovations. While analytically convenient, this assumption is at odds with theoretical models and empirical evidence in the networks literature showing that capabilities vary substantially across firms (magerman2016heterogeneous; chen2023capability; milan2024learningS; acemoglu2025macroeconomics). [^1] Further empirical studies show that the return on R&D is driven by unobserved heterogeneity in innovation efficiency (lentz2008empirical) and organizational competence (henderson1994measuring), rather than spending intensity alone. Indeed, the output elasticity of R&D is highly firm-specific (knott2008r) and depends on the interplay between a firm’s size and innovation strategy (akcigit2018growth).
@@
This paper contributes to three strands of the literature on R&D networks. First, it extends models of endogenous R&D network formation by incorporating firm-level heterogeneity in R&D productivities. Second, it shows how such heterogeneity leads to endogenous asymmetric benefits for connecting firms, which alter firms’ incentives to collaborate and consequently generate a richer set of equilibrium network structures compared to the homogeneous benchmark. Third, it highlights a novel crowding-out effect: while increasing the fraction of high-productive firms in the economy initially raises welfare, beyond a certain threshold a larger fraction of high-productive firms reduces individual R&D efforts in stable network structures, dampening the overall welfare gains from R&D activities.
@@
Our model generalizes the framework introduced in goyal2001r by allowing for heterogeneous firm capabilities to transform R&D efforts into cost reductions. In particular, each firm exerts costly R&D effort [^2] Subject to decreasing marginal returns, in line with the evidence in bloom2020ideas. that lowers the marginal cost of production, and firms form bilateral R&D collaborations [^3] We assume that pairwise link formation is costless, as in goyal2001r and zirulia2012role. Other work that considers linking costs includes goyal2003networks; galeotti2006network; westbrock2010natural; hsieh2024endogenous. with each other to access the benefits of each other’s R&D efforts, while still competing à la Cournot in a homogeneous product market. The introduced heterogeneity in our framework implies that, when firms form bilateral R&D collaborations, the marginal benefit a firm derives from its partner’s effort depends on the partner’s productivity. That is, when a firm with relatively higher productivity increases its effort, the resulting cost reduction for its collaborators is substantial. In contrast, a lower-productivity firm contributes less to its partner’s cost reduction, even if it exerts the same level of R&D effort. Such heterogeneity in firm-level R&D productivities introduces asymmetries in equilibrium R&D efforts and the firm-level gains from R&D collaborations, thus unlocking new results on pairwise stable R&D networks and new transmission channels of these efforts into firm profits and social welfare.
@@
Under heterogeneity, there are two distinct effects of linking with another firm: (i) the productivity effect, and (ii) the connectivity effect — firm-level outcomes depend not only on productivity differences but also on how these firms are linked to other firms in the R&D network. To isolate the effect of heterogeneous productivities from the R&D network effect, we first derive comparative statics on equilibrium efforts and profits for two firms with distinct productivities that occupy symmetric network positions, meaning they face the same network externalities. We show that when these firms are connected, the low-productivity firm experiences a relatively larger increase in profit than the high-productivity firm while exerting a lower R&D effort and therefore bearing a lower R&D cost. Consequently, among two firms with symmetric network positions, whenever the high-productivity firm benefits from forming the link, the low-productivity firm also benefits, showing an ordering of asymmetric gains from linking in favour of the low-productivity firm. Furthermore, our simulation-based findings show that a low-productivity firm always benefits from connecting to a high-productivity firm—while the reverse does not necessarily hold—implying positive/negative benefits from such link formation for low-/high-productivity firms, showing that asymmetric returns go beyond level effects and create differences in link formation incentives. To study the equilibrium network configurations that arise from such asymmetries, we characterize these structures under the widely used equilibrium concept of pairwise stability introduced by jackson1996strategic, as also used in closely related R&D network models, e.g., goyal2001r; goyal2003networks; zirulia2012role; billand2019firm.
@@
Our next result shows that the complete network—which is pairwise stable in the homogeneous framework—ceases to be stable once the productivity gap between firms becomes sufficiently large. Extending stability results to network structures beyond the complete network is known to be intractable due to network externalities, even under the homogeneous benchmark as discussed in goyal2001r, and this intractability is further compounded under heterogeneity. We therefore adopt a simulation-based approach to extract insights on stable R&D networks. Our first key finding in this part shows that the benefit of linking with another firm is monotonically increasing in that other firm’s R&D productivity, and that there exists a productivity threshold for a given firm such that it does not form bilateral links with another firm that has a productivity level below that threshold.
@@
In the rest of the analysis, for simplicity of exposition, we focus on the simplest form of heterogeneity by restricting our setting to two firm types: high- and low-productivity firms, having productivity levels equal to $1$ (normalized) and $\theta<1$, respectively. By using simulations, we then characterize the full set of pairwise stable networks for a given number of firms and type distribution, showing that, under heterogeneity, the set of stable configurations extends beyond the complete network. In particular, while the complete network continues to be stable when the productivity gap between high- and low-type firms is sufficiently small, if this gap is sufficiently large, the class of Positive Assortative (PA) networks becomes stable. These networks are clustered by productivity type, such that firms form bilateral links with all firms of the same type only. Specifically, we establish the existence of a lower threshold productivity $\theta$, denoted by $\underline{\theta}$, above which the complete network is stable, and an upper threshold, $\overline{\theta}$, below which the PA network is stable. This implies that in cases where $\underline{\theta}<\overline{\theta}$, which holds when the fraction of high-productivity firms is not extremely large, an intermediate range for $\theta$ exists for which both networks are pairwise stable. Furthermore, the PA/complete network is the unique stable structure for sufficiently large/small productivity gaps.
@@
Having fully characterized the set of pairwise stable networks, the remainder of the paper shifts attention to a broader implication of heterogeneity beyond firm-level outcomes—its impact on aggregate welfare, measured by the sum of consumer and producer surplus. Simulation-based analysis uncovers a counterintuitive finding: a higher average R&D productivity in the economy —through a higher productivity of low-type firms ($\theta$) or a higher fraction of high-type firms ($\rho$)— may exhibit lower welfare due to (i) the formation of alternative stable R&D network structures and/or (ii) a crowding-out effect of high-productivity firms. First, we find that the PA structure delivers higher welfare compared to the complete network. Consequently, comparing two economies that differ in their productivity gaps, the economy with a smaller gap (where the complete network is stable) exhibits lower welfare than the economy with a larger gap (where the PA network is stable). Second, we find that welfare under PA formation follows an inverted U-shape in the fraction of high-productivity firms: welfare increases with this fraction at low levels, peaks at an intermediate fraction, then declines at high fractions, highlighting a crowding-out effect. Thus, comparing two economies where the PA structure is stable (both having sufficiently large productivity gaps), the economy with the higher fraction of high-productivity firms exhibits lower welfare.
These counterintuitive findings contradict the conventional wisdom in economic models that increasing productivity in an economy is welfare-enhancing. When collaborations between firms and the resulting interactions in the overall economy are taken into account, higher average R&D productivity need not be welfare-enhancing if it alters stable network structures and their associated network effects. A relevant policy implication is that innovation and productivity-enhancing policies should account for their impact on endogenous R&D collaborations and effort, as endogenous responses in the network structure may offset or even reverse the intended welfare gains. Moreover, an avenue for future research is to empirically analyze these findings from a cross-country perspective or by examining changes in innovation productivities and collaborations between firms over time within a given region. Such analysis can offer an alternative explanation for the productivity puzzle observed in developed economies recently—while firms may become more productive in innovation at the individual level, this does not necessarily translate into higher aggregate welfare (or observed overall productivity) due to altered incentives in the innovation landscape.
Our approach differs from existing contributions by allowing for endogenous asymmetric benefits to link formation arising from productivity differentials. For example, zirulia2012role introduces pre-defined partner-specific spillovers, highlighting how firms with unique technological capabilities can become central hubs in their markets. In a similar direction, billand2019firm introduces connection-specific benefits from linking, where link-specific benefits are the same for a linked pair but can differ across the type of pairing. This existing work exhibits distinct mechanisms and results from ours. More specifically, in these models, the marginal cost reduction of pairwise linking is the same for two connecting firms, regardless of any underlying heterogeneity between them. In contrast to these models, our framework allows for such marginal cost reductions to be endogenously asymmetric, generating different network formation incentives and pairwise stable structures with distinct welfare implications. Additionally, while some other models explore the dynamic formation of the R&D network and evolution of technology diffusion (such as bischi2012dynamicP1; hsieh2024endogenous), our model focuses on a static environment that allows us to fully characterize the implications of heterogeneity for link formation, effort allocation, and aggregate welfare.
@@
The remainder of the paper is structured as follows. Section 2 outlines the theoretical framework and the multi-stage game setup, providing our main theoretical findings. By simplifying to a two-type productivity setup, Section 3 provides a full characterization of pairwise stable networks for $n=4$ firms, and extends insights on firm-level outcomes, pairwise stability and welfare through simulations. Section 4 provides our closing remarks and discusses how our model can have implications for industrial policy.
@@ 2 Model
There exists a set of firms $\mathcal{N}=\{1,\ldots,n\}$ competing in a single-product oligopolistic market and we employ a complete information multi-stage game setup similar to goyal2001r. The timing of actions is as follows:
@@
    Network Formation. Firms engage in bilateral R&D alliances with one another, represented as an undirected link between two collaborators. Following jackson1996strategic, we use the standard notion of pairwise stability as the equilibrium concept. Letting $\mathcal{G}^{n}$ denote the set of all possible binary $n\times n$ adjacency matrices, an R&D network is represented by an zero-diagonal, symmetric adjacency matrix $\mathbf{G}\in\mathcal{G}^{n}$ in which $G_{ij}$ equals 1 if there is a collaboration between firms $i$ and $j$, and 0 otherwise.
@@
    R&D Investment. Given the network $\mathbf{G}$, firms choose their level of R&D effort ($\tilde{e}_{i}$) with a quadratic cost $\text{C}(\tilde{e}_{i})=\phi_{i}\cdot\tilde{e}_{i}^{2}$, where $\phi_{i}>0$ is the firm specific steepness of the cost function. Then, firms share their R&D efforts with their collaborators in the network. As a result, each firm’s marginal cost of production ($c_{i}$) is:
@@
We depart from the literature by adding heterogeneity in R&D productivities into the baseline model described above as introduced by goyal2001r. Specifically, while earlier studies assume that $\phi_{i}=\phi$ for all firms, in our framework $\phi_{i}$ captures the inverse of firm $i$’s R&D productivity. A higher value of $\phi_{i}$ implies the firm must incur a greater R&D cost to achieve a given efficiency (or technology) level as measured by its marginal production cost $c_{i}$.
For ease of interpretation, we normalize R&D costs relative to the most R&D-efficient firm in the market. Let $\phi$ denote the cost coefficient of this most productive firm, defined as $\phi\coloneqq\min\{\phi_{1},\ldots,\phi_{n}\}.$ We then define the relative R&D productivity of firm $i$ with respect to the most productive firm as $\theta_{i}\coloneqq\sqrt{\phi/\phi_{i}}$ and a rescaled R&D effort variable as $e_{i}\coloneqq\tilde{e}_{i}/\theta_{i}$. Given this normalization, $\theta_{i}\in(0,1]$ and the most productive firm(s) having $\max_{i\in\mathcal{N}}\theta_{i}=1$.

### v2

# The R&D Productivity Puzzle: Innovation Networks with Heterogeneous Firms

@@ Abstract
We introduce heterogeneous R&D productivities into an endogenous R&D network formation model, generalizing the framework of Goyal and Moraga-González (2001). Heterogeneous productivities endogenously create asymmetric gains from collaboration: less productive firms benefit disproportionately from links, while more productive firms exert greater R&D effort and incur higher costs. When productivity gaps are sufficiently large, more productive firms experience lower profits from collaborating with less productive partners. As a result, the complete network—stable under homogeneity—becomes unstable, and the positive assortative (PA) network, in which firms cluster by R&D productivity, emerges as pairwise stable. Using simulations, we show that the clustered structure delivers higher welfare than the complete network; nevertheless, welfare under this formation follows an inverted U-shape as the fraction of high-productivity firms increases, reflecting crowding-out effects at high fractions. Altogether, we uncover an R&D productivity puzzle: economies with higher average R&D productivity may exhibit lower welfare through (i) the formation of alternative stable networks, or (ii) a crowding-out effect of high-productivity firms. Our findings highlight that productivity is a structural force reshaping the organization of innovation. Productivity-enhancing policies must therefore account for their effects on endogenous R&D alliances and effort, as they may reverse the intended welfare gains. JEL Classification: D21, D85, O30.
@@ 1 Introduction
Innovation is widely recognized as a fundamental driver of economic growth in modern societies (Romer, 1990; Jones, 1995). A key engine of the innovation process is private firms’ investment in R&D (Grossman and Helpman, 1991; Aghion and Howitt, 1992), a substantial share of which takes place through collaborative R&D alliances, as documented empirically by Cassiman and Veugelers (2002); Belderbos et al. (2004); Calero et al. (2007).
@@
In this paper, we study R&D alliances and their implications for firms’ profits and aggregate welfare in an R&D network formation model with heterogeneous R&D productivities. A growing theoretical literature, including Goyal and Moraga-González (2001); Goyal and Joshi (2003); König et al. (2019); Hsieh et al. (2025), studies R&D networks where firms improve their R&D efficiencies by forming collaborations while still competing with each other in the product market. However, these models typically assume that firms are homogeneous in their ability to transform R&D effort into cost-reducing innovations. While analytically convenient, this assumption is at odds with theoretical models and empirical evidence showing that capabilities vary substantially across firms (Magerman et al., 2016; Chen et al., 2023; Milán, 2019; Acemoglu and Tahbaz-Salehi, 2025). [^1] Further empirical studies show that the return on R&D is driven by unobserved heterogeneity in innovation efficiency (Lentz and Mortensen, 2008) and organizational competence (Henderson and Cockburn, 1994), rather than spending intensity alone. Indeed, the output elasticity of R&D is highly firm-specific (Knott, 2008) and depends on the interplay between a firm’s size and innovation strategy (Akcigit and Kerr, 2018). We show that once this assumption is relaxed and firms are heterogeneous, a productivity puzzle is introduced: economies with larger overall R&D productivity can exhibit lower welfare.
@@
This paper contributes to three strands of the literature on R&D networks. First, it extends models of endogenous R&D network formation by incorporating firm-level heterogeneity in R&D productivities. Second, this methodological approach demonstrates how heterogeneity leads to endogenous asymmetric benefits for connecting firms, which alter firms’ incentives to collaborate and consequently generate a richer set of equilibrium network structures compared to the homogeneous benchmark. Third, we introduce and explain the mechanisms underlying the counterintuitive R&D productivity puzzle: economies with higher average R&D productivity may exhibit lower aggregate welfare through (i) the formation of denser, stable R&D networks with lower individual efforts or (ii) a crowding-out effect of high-productivity firms. Welfare losses arise since the connectivity effect dominates the productivity effect in determining welfare.
@@
Our model generalizes the framework introduced in Goyal and Moraga-González (2001) by allowing for heterogeneous firm capabilities to transform R&D efforts into cost reductions. In particular, each firm exerts costly R&D effort [^2] Subject to decreasing marginal returns, in line with the evidence in Bloom et al. (2020).—with heterogeneous returns—to lower the marginal cost of production. Firms also form bilateral R&D collaborations [^3] We assume that pairwise link formation is costless, as in Goyal and Moraga-González (2001) and Zirulia (2012). Other work that considers linking costs includes Goyal and Joshi (2003); Galeotti et al. (2006); Westbrock (2010); Hsieh et al. (2025). with each other to access the benefits of each other’s R&D efforts, while still competing à la Cournot in a homogeneous product market. This generalization to firm-level heterogeneity implies that, when firms form bilateral R&D collaborations, the marginal benefit a firm derives from its partner’s effort depends on the partner’s productivity. That is, when a firm with relatively higher productivity increases its effort, the resulting cost reduction for its collaborators is substantial. In contrast, a lower-productivity firm contributes less to its partner’s cost reduction, even if it exerts the same level of R&D effort. Such heterogeneity introduces asymmetries in equilibrium R&D efforts and the firm-level gains from R&D collaborations, thus unlocking new results on pairwise stable R&D networks and new transmission channels of these efforts into firm profits and social welfare.
@@
Under heterogeneity, there are two distinct effects of linking with another firm: (i) the productivity effect, and (ii) the connectivity effect—firm-level outcomes depend on both productivity differences and network externalities arising from R&D alliances. To isolate the former effect, we first derive comparative statics on equilibrium efforts and profits for two firms with distinct productivities that occupy symmetric network positions, meaning they face the same network externalities. Our first result shows that when such two firms are connected, the low-productivity firm experiences a relatively larger increase in profit than the high-productivity firm while exerting a lower R&D effort and therefore bearing a lower R&D cost. Consequently, among two firms with symmetric network positions, whenever the high-productivity firm benefits from forming the link, the low-productivity firm also benefits, showing an ordering of asymmetric gains from linking in favor of the low-productivity firm. Furthermore, our simulation-based findings show that such asymmetric returns go beyond level effects, and in fact, create differences in link formation incentives: a low-productivity firm always benefits from connecting to a high-productivity firm, while the reverse does not necessarily hold. This implies there are positive/negative benefits from link formation for low-/high-productivity firms.
@@
Building on these fundamental insights, we study equilibrium network configurations under the widely adopted concept of pairwise stability introduced by Jackson and Wolinsky (1996) and employed in closely related R&D network models, e.g., Goyal and Moraga-González (2001); Goyal and Joshi (2003); Zirulia (2012); Billand et al. (2019). Our next theoretical result shows that the complete network—which is pairwise stable in the homogeneous framework—ceases to be stable once the productivity gap between firms becomes sufficiently large. The reason is that the benefit of linking with another firm is monotonically increasing in that other firm’s R&D productivity, which means there exists a productivity threshold for a given firm such that it does not form bilateral links with any firm that has a productivity level below that threshold.
Extending stability results to network structures beyond the complete network is known to be intractable due to network externalities, even under the homogeneous benchmark as discussed in Goyal and Moraga-González (2001). This intractability is further compounded under heterogeneity. We therefore adopt a simulation-based approach to extract insights on the set of stable R&D networks and their welfare implications.
@@
In the rest of the analysis, for simplicity of exposition, we focus on the simplest form of heterogeneity by restricting our setting to two firm types: high- and low-productivity firms, with productivity levels normalized to $1$ and $\theta\in(0,1]$, respectively. By using simulations, we then characterize the full set of pairwise stable networks for a given number of firms and type distribution, showing that, under heterogeneity, the set of stable configurations extends beyond the complete network. In particular, while the complete network continues to be stable when the productivity gap between high- and low-type firms is sufficiently small, if this gap is sufficiently large (low $\theta$), the class of Positive Assortative (PA) networks becomes stable. These networks are clustered by productivity type, such that firms form bilateral links with all firms of the same type only. Specifically, we establish the existence of a lower threshold productivity $\theta$, denoted by $\underline{\theta}$, above which the complete network is stable, and an upper threshold, $\overline{\theta}$, below which the PA network is stable. This implies that in cases where $\underline{\theta}<\overline{\theta}$, which holds when the fraction of high-productivity firms is not extremely large, an intermediate range for $\theta$ exists for which both networks are pairwise stable. Furthermore, the PA/complete network is the unique stable structure for sufficiently large/small productivity gaps.
@@
The remainder of the paper shifts attention beyond firm-level outcomes to a broader implication of heterogeneity: aggregate welfare as measured by the sum of consumer and producer surplus. Simulation-based analysis uncovers the R&D productivity puzzle: a higher average R&D productivity in the economy —achieved through a higher productivity of low-type firms ($\theta$) or a higher fraction of high-type firms ($\rho$)— may exhibit lower welfare. This happens due to (i) the formation of alternative stable R&D network structures and/or (ii) a crowding-out effect of high-productivity firms. First, we find that the clustered structure (PA) delivers higher welfare compared to the complete network. Consequently, comparing two economies that differ in their productivity gaps, the economy with a smaller gap (in which the complete network is stable) exhibits lower welfare than the economy with a larger gap (in which the PA network is stable). Second, we find that welfare under PA formation follows an inverted U-shape in the fraction of high-productivity firms: welfare increases with this fraction at low levels, peaks at an intermediate fraction, then declines at high fractions, highlighting a crowding-out effect.
These counterintuitive findings challenge the conventional economic wisdom that higher productivity is welfare-enhancing (Romer, 1990; Jones, 1995; Jones and Williams, 2000; Basu et al., 2022). Viewed through the lens of R&D networks, we show that endogenous connectivity incentives can alter the welfare effects of R&D productivity improvements. At a deeper level, our findings reveal that productivity affects both the efficiency and organization of innovation across firms. An important policy implication is that innovation-enhancing policies should account for their impact on R&D alliances, as changes in stable network structures may offset or even reverse intended welfare gains. A natural avenue for future research is to test these mechanisms empirically, either in a cross-country setting or by examining changes in firms’ innovation productivity and collaboration patterns over time within a region.
Our approach differs from existing contributions by allowing endogenous asymmetric benefits from link formation to arise from R&D productivity gaps between firms. Zirulia (2012) introduces exogenous, partner-specific spillovers, showing how firms with unique technological capabilities can become central hubs in their markets. Similarly, Billand et al. (2019) studies connection-specific benefits, where the gains from linking are symmetric within a given pair but differ across types of pairings. While these models generate rich network structures, the marginal cost reduction from forming a link is symmetric for the two connected firms, regardless of underlying heterogeneity.
In contrast, our framework allows marginal cost reductions from collaboration to be endogenously asymmetric, generating fundamentally different incentives for link formation and leading to pairwise stable network structures with distinct welfare implications. Finally, whereas some related work studies the dynamic formation of R&D networks and the evolution of technology diffusion (e.g., Bischi and Lamantia, 2012; Hsieh et al., 2025), our model focuses on a static environment, which allows us to analyze how heterogeneity shapes link formation, R&D effort allocation, and aggregate welfare.
@@
The remainder of the paper is as follows. Section 2 outlines the theoretical framework and the multi-stage game setup, providing our main theoretical findings. By simplifying to a two-type productivity setup, Section 3 provides the full set of pairwise stable networks for $n=4$ firms, and extends insights on firm-level outcomes, pairwise stability and welfare through simulations. Section 4 provides our closing remarks and discusses how our model can have implications for industrial policy.
@@ 2 Model
There exists a set of firms $\mathcal{N}\coloneqq\{1,\ldots,n\}$ competing in a single-product oligopolistic market and we employ a complete information multi-stage game setup similar to Goyal and Moraga-González (2001). The timing of actions is as follows:
@@
    Network Formation. Firms engage in bilateral R&D alliances with one another, represented as an undirected link between two collaborators. Following Jackson and Wolinsky (1996), we use the standard notion of pairwise stability as the equilibrium concept. Letting $\mathcal{G}^{n}$ denote the set of all possible binary $n\times n$ adjacency matrices, an R&D network is represented by a zero-diagonal, symmetric adjacency matrix $\mathbf{G}\in\mathcal{G}^{n}$ in which $G_{ij}$ equals 1 if there is a collaboration between firms $i$ and $j$, and 0 otherwise.
@@
    R&D Investment. Given the network $\mathbf{G}$, any firm $i$ chooses their level of R&D effort ($\tilde{e}_{i}$) facing a quadratic cost $\text{C}(\tilde{e}_{i})=\phi_{i}\cdot\tilde{e}_{i}^{2}$, where $\phi_{i}>0$ is the firm-specific cost function steepness. Firms then share their R&D efforts with their collaborators in the network. As a result, each firm’s marginal cost of production ($c_{i}$) is:
@@
We depart from the literature by adding heterogeneity in R&D productivities into the baseline model described above as introduced by Goyal and Moraga-González (2001). Specifically, while earlier studies assume that $\phi_{i}=\phi$ for all firms, in our framework $\phi_{i}$ captures the inverse of firm $i$’s R&D productivity. A higher value of $\phi_{i}$ implies the firm must incur a greater R&D cost to achieve a given efficiency (or technology) level as measured by its marginal production cost $c_{i}$.
For ease of interpretation, we normalize R&D costs relative to the most R&D-efficient firm in the market. Let $\phi$ denote the cost coefficient of this most productive firm, defined as $\phi\coloneqq\min\{\phi_{1},\ldots,\phi_{n}\}.$ We then define the relative R&D productivity of firm $i$ with respect to the most productive firm as $\theta_{i}\coloneqq\sqrt{\phi/\phi_{i}}$ and a rescaled R&D effort variable as $e_{i}\coloneqq\tilde{e}_{i}/\theta_{i}$. Given this normalization, $\theta_{i}\in(0,1]$ and the most productive firm(s) have $\max_{i\in\mathcal{N}}\theta_{i}=1$.